OpenStudy (anonymous):

The equation of a circle is (x + 3)^2 + (y + 5)^2 = 8. Determine the coordinates of the center of the circle.

hartnn (hartnn):

for \(\Large (x-h)^2+(y-k)^2 =r^2 \) the center is h,k so how about \(\Large (x+3)^2+(y+5)^2 =8 \)

hartnn (hartnn):

compare those 2 equations

OpenStudy (ahsome):

The x coordinate is the -h value, while the y coordinate is the -k value. \[(x+3)^2+(y+5)^2=8\] In this case, the h value is 3 and the k value is 5 Since the x-coordinate is -h: \[x=-h\] \[x=-1*3\] \[x=-3\] This is the same with the y-coordinate \[y=-k\] \[y=-1*5\] \[y=-5\]

OpenStudy (anonymous):

I got (3,5)

OpenStudy (anonymous):

I was late.

OpenStudy (ahsome):

Remember, the coordinates are the NEGATIVE of the h, k value. See above